OpenStudy (anonymous):

x''=4y+e^(t) y''=4x-e^(t) Solve the differential equations by systematic elimination

OpenStudy (perl):

so these are parametric equations

OpenStudy (anonymous):

\[\begin{cases}x''(t)=4y+e^t\\y''=4x-e^t\end{cases}\] Solving for \(x\) in the second equation gives \[x=\frac{y''-e^t}{4}\] Differentiating twice, you have \[x''=\frac{y^{(4)}-e^t}{4}\] Substitute into the first equation: \[\begin{align*}\frac{y^{(4)}-e^t}{4}&=4y+e^t\\\\ \frac{1}{4}y^{(4)}-4y&=\frac{5}{4}e^t \end{align*}\] The characteristic equation is \[\frac{1}{4}r^4-4=0\] which has four roots. \[r^4-16=(r-2)(r+2)(r-2i)(r+2i)=0~~\implies~~r=\pm2,\pm2i\] so your homogeneous solution is \[y_h=C_1e^{2t}+C_2e^{-2t}+C_3\cos2t+C_4\sin2t\] For the nonhomogeneous part, you can use the method of undetermined coefficients. As your guess function, use \[y=Ae^t~~\implies~~y^{(n)}=Ae^t\] \[\begin{align*}\frac{A}{4}e^t-4Ae^t&=\frac{5}{4}e^t\\\\ -\frac{15}{4}Ae^t&=\frac{5}{4}e^t\\\\ A&=-\frac{1}{3} \end{align*}\] and hence the general solution is \[y(t)=y_h-\frac{1}{3}e^t\]

OpenStudy (perl):

i am confused by the nonhomogenous part, how do you 'guess' the right function

OpenStudy (anonymous):

It's the standard procedure of the method of undetermined coefficients. If you have a second order ODE with constant coefficients, i.e. of the form \[ay''+by'+cy=f(x)\] you can check to see if a variant of \(f(x)\) fits as a solution for \(y\). As a particular example, suppose \(f(x)\) is a polynomial. Then we can suppose \(y_p\), our guess solution, is also a polynomial, because derivatives of polynomials are also polynomials. This method generally works if \(f(x)\) is a linear combination of sines, cosines, exponentials, and (I think, I could be wrong) what I might call "polynomial-periodic" functions, like \(x^2\sin2x\).